# A demonstration of the  class file for SAGE Publications

### v1

@@ 4.1 Semantic Keypoint Discovery
To discover semantic keypoints, we first extract the contour and skeleton of clothes, which serve as an effective abstraction, capturing the geometric structure of clothes and encoding the human body. Specifically, we employ OWL-V2 (Minderer et al., 2024) to generate the bounding box of clothes on the prototype image based on the text prompt. Later, the bounding box will serve as the box prompts for the Segment-Anything Model (SAM) (Kirillov et al., 2023) to generate the corresponding masked image $I_{m}$ of the clothes. We then extract the contour of clothes from the masked image $I_{m}$ and perform skeletonization by calculating equidistance to the boundaries (Lee et al., 1994), resulting in the skeleton. To get keypoints from the skeleton and boundary, we then apply farthest point sampling (fps) (Eldar et al., 1997) on the boundary and skeleton to sample candidate keypoints $\boldsymbol{K}_{fps}$. Farthest point sampling is an effective method for obtaining diverse, spatially distributed candidate keypoints that are geometrically salient.
After obtaining candidate keypoints $\boldsymbol{K}_{fps}$, selecting semantic keypoints requires semantic understanding of clothes and clothes manipulation. Therefore, we leverage the commonsense knowledge of the VLM and employ mark-based visual prompting (Fang et al., 2024a) to select semantic keypoints and describe them with natural language. In this paper, we use GPT-4o as the VLM. The prompt is multi-modal and consists of two parts: (i) text prompt, and (ii) marked image. In the text prompt, we provide context information to the VLM, including the definition of semantic keypoints, the expected behavior of the VLM, and the desired response format. In the marked image, we draw visual marks on candidate keypoints $\boldsymbol{K}_{fps}$ and label them with numbers. Through mark-based visual prompting, we can get the semantic keypoints $\boldsymbol{K}_{sem}=\{p_{i},s_{i}\}_{i=1}^{n}$, where $p_{i}$, $s_{i}$ refer to the pixel position and language description of $i$-th semantic keypoint, respectively, and $n$ represents the total number of semantic keypoints.
@@ 4.2 Semantic Keypoint Matching
Given an RGB observation image of novel clothes, we first retrieve the most relevant prototype image by computing the Best Buddies Similarity (BBS) using features extracted from the vision foundation model DINOv2 (Oquab et al., 2023). BBS is a robust method for measuring structural similarity between two images. It identifies pairs of local patches—one from the observation image and one from the prototype image—that are mutual nearest neighbors in the feature space. These pairs, referred to as best buddy pairs, capture strong local correspondences. The BBS score is defined as the proportion of local patches in the observation image that form such pairs. By evaluating the BBS score across all prototype images, we retrieve the one that is most structurally similar to the observation image. To enhance robustness, as shown in Figure 4, we apply a set of rotations to the observation image and select the rotation that yields the highest BBS score.
@@
After retrieving the most relevant prototype image $I_{r}$ annotated with semantic keypoints, we then perform a coarse-to-fine semantic keypoint matching on novel clothes. Given the novel clothes’ RGB observation image $I_{q}$ and a semantic keypoint position $p_{i}$ on the prototype image, we first match $p_{i}$ to a coarse region $\Omega$ on the observation image $I_{q}$. This stage is crucial for capturing global context and avoiding local minima. Specifically, we employ VLM with mark-based visual prompting for coarse region proposals. We first mark the observation image $I_{q}$ with an overlaid grid, where each grid cell is indexed by a unique text label, and mark the prototype image $I_{r}$ with a dot on $p_{i}$. We then prompt the VLM with the marked observation and prototype image to select the grid indices corresponding to the prototype semantic keypoint position $p_{i}$. To get precise semantic keypoint positions, we then perform fine-grained keypoint matching on the region $\Omega$ with vision foundation models DINOv2 (Oquab et al., 2023) and stable diffusion (Rombach et al., 2022). Borrowing the method from (Zhang et al., 2024a), we feed $I_{r}$ and $I_{q}$ into vision foundation models, respectively, and get feature maps $\mathcal{F}_{r}$ and $\mathcal{F}_{q}$. In this way, we can get the feature $\mathcal{F}_{q}(p)$ of each point $p$ on the region $\Omega$ and the prototype semantic keypoint’s feature $\mathcal{F}_{r}(p_{i})$. We then calculate the similarity score between them to achieve fine-grained keypoint matching:
Figure 6: Skill Library. Our skill library consists of 5 basic skills: grasp, moveto, release, rotate and pull.
$p_{i}^{\prime}=\arg\max_{p\in\Omega}\langle\mathcal{F}_{q}(p),\mathcal{F}_{r}(p_{i})\rangle$ (1)
[…]
@@ 5.1 Skill Library
For general-purpose clothes manipulation, we build a skill library at first. To ensure that the skill library is general and sufficient for various clothes and manipulation tasks, we first prompt the LLM to propose basic skills of clothes manipulation with the chain of thought prompting (Wei et al., 2022). Specifically, the LLM is prompted to (i) generate examples of clothes manipulation tasks, (ii) break down these tasks into basic actions, and (iii) summarize the actions identified in step (iii) to derive a set of basic skills. In this way, we identify the basic skill set $\mathcal{A}$, including grasp: pick up a part of the clothes, moveto: transport the clothes to a target position, release: drop a part of the clothes, rotate: rotate the clothes, and pull: stretch the clothes for flattening. Figure 6 visualizes the basic skills discovered by the LLM. Since semantic keypoints serve as the effective affordances for clothes manipulation and define where clothes are more likely to be manipulated, we implement these basic skills as policies parameterized by contact point positions. Specifically, we define heuristic rules to create waypoints based on the contact point positions for each basic skill. Once the waypoints are determined, a motion planning algorithm is applied to generate the complete trajectory for the robot’s execution. Specifically, we implement each basic skill as follows:
@@
release()()( ): The skill release drops clothes parts by opening grippers. After the release action, the observation is updated, and CLASP extracts new semantic keypoints.
@@
[…]
@@ 5.2 Task Planning
Figure 7: VLM-powered Task Planner. We implement a closed-loop VLM-powered task planner capable of dynamic replanning based on observations, overall task completion checking, and plan verification.
After establishing the skill library for clothes manipulation, we employ a VLM for task planning based on semantic keypoints, taking advantage of its extensive commonsense knowledge (Zhao et al., 2024) gained from internet-scale data. To generate the task plan, we leverage the multi-modal reasoning capabilities of the VLM and employ visual prompting (Figure 7). Specifically, the prompt includes a meta-prompt and few-shot examples. The meta-prompt defines the task objective, describes the available basic skills, and specifies the input-output format. To improve the robustness of task planning, we leverage in-context learning by providing the VLM with few-shot examples consisting of language instructions, an observation image annotated with extracted semantic keypoints, and the corresponding task plan. These examples not only illustrate the expected input-output format but also demonstrate how to utilize the basic skills from the predefined skill library.
Algorithm 1 Task Planning and Verification
0: Language instruction $\mathcal{L}$, observation image $I_{q}$, semantic keypoints $\boldsymbol{K}_{sem}^{\prime}$, skill set $\mathcal{A}$, max attempts $N_{\max}$
0: Verified task plan or failure reason
1:Step 1: Task Completion check
[…]
3:  return  “Task Completed”
4:end if
[…]
@@
7:  $n_{\text{plan}}\leftarrow n_{\text{plan}}+1$
8:  Step 2: Task Plan generation
[…]
10:  Step 3: Plan Verification
11:  failure_reason $\leftarrow$ None
12:  for $\boldsymbol{w}_{j}=(a_{j},c_{j}^{1},c_{j}^{2})$ in $\{\boldsymbol{w}_{j}\}$ do
13:   if $a_{j}\notin\mathcal{A}$ or $c_{j}^{1,2}\notin\boldsymbol{K}_{sem}^{\prime}$ then
14:    failure_reason $\leftarrow$ “Invalid subtask $\boldsymbol{w}_{j}$
15:    break
[…]
17:    failure_reason $\leftarrow$ “Infeasible motion in $\boldsymbol{w}_{j}$
18:    break
19:   end if
20:  end for
21:until failure_reason = None or $n_{\text{plan}}=N_{\max}$
22:if failure_reason = None then
23:  return  $\{\boldsymbol{w}_{j}\}_{j=1}^{H}$
24:else
25:  return  failure_reason
26:end if
The task planning procedure is described in Algorithm LABEL:alg:task_planning. Given a language instruction $\mathcal{L}$, observation image $I_{q}$, and extracted semantic keypoints $\boldsymbol{K}_{sem}^{\prime}$, the VLM decides whether the task is completed and generates task plan $\{\boldsymbol{w}_{j}\}_{j=1}^{H}$, where $H$ denotes the number of subtasks. Each subtask consists of a basic skill and two contact points for the dual-arm manipulation system: $\boldsymbol{w}_{j}=(a_{j},c_{j}^{1},c_{j}^{2})$. The basic skill $a_{j}$ is selected from the predefined set $\mathcal{A}$ introduced in Sec 5.1. The contact points $c_{j}^{1}$ and $c_{j}^{2}$ are selected from extracted semantic keypoints $\boldsymbol{K}_{sem}^{\prime}$. Semantic keypoints define where clothes can be manipulated and serve as contact point candidates in high-level task planning. It should be noted that both $c_{j}^{1}$ and $c_{j}^{2}$ are optional, depending on the number of required contact points. For example, no contact point is required for the skill release. After obtaining the task plan, CLASP proceeds with plan verification by sequentially validating each subtask. For each subtask, CLASP first performs syntax and semantic analysis to ensure that the invoked skill is defined within the skill library and contact points correspond to valid semantic keypoints. It then verifies the subtask at the motion level. Specifically, CLASP invokes a policy from the skill library to create waypoints and perform motion planning to generate the trajectory. The subtask is verified if it passes motion planning. If a subtask fails verification, it is reported as the failure reason to the VLM, prompting the generation of a new task plan.
@@
After obtaining the verified task plan, CLASP executes the subtasks sequentially. Following the completion of each subtask, the RGB-D observation is updated. However, due to occlusion caused by the robot, semantic keypoints are only extracted and updated after the release action. Once the semantic keypoints are updated, the VLM is prompted to replan based on the new observation and updated keypoints. The above process repeats until the task is completed.
@@ 6 Experiments
Figure 8: Qualitative Results of Semantic Keypoint Extraction. Our method can achieve open-category semantic keypoint extraction under irregular deformation and occlusion.
@@
To answer this, we evaluate the performance of CLASP’s submodules—semantic keypoint extraction and task planning. Results in Sec. 6.1 show that our method can extract effective semantic keypoints on a wide range of clothes even under occlusion and deformation. Results in Sec. 6.2 show that our task planning method can generate reliable and executable task plans.
@@ 6.1 Semantic Keypoint Extraction Experiments
Table 1: Semantic Keypoint Extraction Results.
@@
We first evaluated the performance of CLASP in extracting effective semantic keypoints. Since no groundtruth semantic keypoints are available, we provide annotated prototype images to three human experts and ask them to label semantic keypoints on novel clothes as ground truth. In this way, we establish a testing set of 120 images from 15 different clothes. We then compare the semantic keypoints extracted by CLASP against those labeled by the experts. In these images, clothes have diverse configurations, poses, occlusions, and deformations.
The evaluation utilizes two standard metrics for keypoint extraction: Average Keypoint Distance (AKD) and Average Precision (AP). AKD calculates the average distance between the extracted keypoints and the groundtruth, while AP measures the proportion of correctly extracted keypoints under different thresholds. Since the resolution of the observed RGB-D images is 720×1280, we used thresholds of 15, 30, and 45 pixels for AP calculation.
@@
Considering the resolution and uncertainty in expert labeling, such as identifying the center of clothes, the results can demonstrate that our proposed method performs well in extracting semantic keypoints. For each clothes type, we only need to provide one prototype image without any annotation. Figure 8 shows some qualitative results of semantic keypoint extraction. The experiment results demonstrate that our two-stage semantic keypoint extraction pipeline is effective in extracting semantic keypoints of clothes even under occlusion and deformation. The prior knowledge from the vision language model and vision foundation models ensures the generalization to diverse clothes types. Besides, coarse-to-fine keypoint matching ensures spatial precision and makes semantic keypoint extraction robust.
Table 2: Task planning evaluation results.

### v2

@@ 4.1 Keypoint Discovery
To discover semantic keypoints, we first extract the contour and skeleton of clothes, which serve as an effective abstraction, capturing the geometric structure of clothes and encoding the human body. Specifically, we employ OWL-V2 (Minderer et al., 2024) to generate the bounding box of clothes on the prototype image based on the text prompt. Later, the bounding box will serve as the box prompts for the Segment-Anything Model (SAM) (Kirillov et al., 2023) to generate the corresponding masked image $I_{m}$ of the clothes. We then extract the contour of clothes from the masked image $I_{m}$ and perform skeletonization by calculating equidistance to the boundaries (Lee et al., 1994), resulting in the skeleton. To get keypoints from the skeleton and boundary, we then apply farthest point sampling (fps) (Eldar et al., 1997) on the boundary and skeleton to sample candidate keypoints $\boldsymbol{K}_{c}$. Farthest point sampling is an effective method for obtaining diverse, spatially distributed candidate keypoints that are geometrically salient.
After obtaining candidate keypoints $\boldsymbol{K}_{c}$, selecting semantic keypoints requires semantic understanding of clothes and clothes manipulation. Therefore, we leverage the commonsense knowledge of the VLM and employ mark-based visual prompting (Fang et al., 2024a) to select semantic keypoints and describe them with natural language. In this paper, we use GPT-4o as the VLM. The prompt is multi-modal and consists of two parts: (i) text prompt, and (ii) marked image. In the text prompt, we provide context information to the VLM, including the definition of semantic keypoints, the expected behavior of the VLM, and the desired response format. In the marked image, we draw visual marks on candidate keypoints $\boldsymbol{K}_{c}$ and label them with numbers. Through mark-based visual prompting, we can get the prototype semantic keypoints $\boldsymbol{K}_{0}=\{p_{i,0},\,s_{i,0}\}_{i=1}^{n}$, where $p_{i,0}$, $s_{i,0}$ denote the pixel position and language description of $i$-th semantic keypoint, respectively, and $n$ is the number of keypoints.
@@ 4.2 Keypoint Matching
The pipeline of semantic keypoint matching is shown in Figure 4. Given an RGB observation image of novel clothes, we first retrieve the most relevant prototype image by computing the Best-Buddies Similarity (BBS) using features extracted from the vision foundation model DINOv2 (Oquab et al., 2023). DINOv2 can extract rich and transferable visual features efficiently. BBS is a robust metric for measuring structural similarity between two images. It identifies pairs of local patches—one from the observation image and one from the prototype image—that are mutual nearest neighbors in the feature space. These pairs, referred to as best buddy pairs, capture strong local correspondences. The BBS score is defined as the proportion of local patches in the observation image that form such pairs. By evaluating the BBS score across all prototype images, we retrieve the one that is most structurally similar to the observation image. To enhance robustness, we apply a set of rotations to the observation image and select the rotation that yields the highest BBS score.
@@
After retrieving the most relevant prototype image $I_{0}$ annotated with semantic keypoints, we then perform a coarse-to-fine semantic keypoint matching on novel clothes. Given the novel clothes’ RGB observation image $I$ and a semantic keypoint position $p_{i,0}$ on the prototype image, we first match $p_{i,0}$ to a coarse region $\Omega$ on the observation image $I$. This stage is crucial for capturing global context and avoiding local minima. Specifically, we employ VLM with mark-based visual prompting for coarse region proposals. We first mark the observation image $I$ with an overlaid grid, where each grid cell is indexed by a unique text label, and mark the prototype image $I_{0}$ with a dot on $p_{i,0}$. We then prompt the VLM with the marked observation and prototype image to select the grid indices corresponding to the prototype semantic keypoint position $p_{i,0}$. To obtain precise semantic keypoint positions, we then perform fine-grained keypoint matching on the region $\Omega$ with vision foundation models. Borrowing the method from (Zhang et al., 2024a), we fuse the feature maps from DINOv2 (Oquab et al., 2023) and Stable Diffusion (SD) (Rombach et al., 2022). The fused feature effectively captures both local and global information, thereby enhancing the robustness of semantic keypoint matching. Specifically, we feed the observation image $I$ and the prototype image $I_{0}$ into the fused feature extractor $\mathcal{F}(\cdot)$, which produces their corresponding fused feature maps $\mathcal{F}(I)$ and $\mathcal{F}(I_{0})$. For each point $p\in\Omega$ on the observation image, we obtain its feature representation $\mathcal{F}(I,\,p)$, and for each prototype semantic keypoint $p_{i,0}$, we obtain its corresponding prototype feature $\mathcal{F}(I_{0},\,p_{i,0})$. We then compute the cosine similarity between $\mathcal{F}(I,\,p)$ and $\mathcal{F}(I_{0},\,p_{i,0})$ to perform fine-grained keypoint matching:
$p_{i}=\arg\max_{p\in\Omega}\langle\mathcal{F}(I,\,p),\,\mathcal{F}(I_{0},\,p_{i,0})\rangle,$ (1)
where $\langle\cdot,\cdot\rangle$ denotes the cosine similarity between two feature vectors. Here, $p_{i}$ represents the pixel position of the matched keypoint on the observation image $I$ corresponding to the prototype keypoint $p_{i,0}$. Repeating this process for all prototype semantic keypoints $\boldsymbol{K}_{0}$, we obtain the matched semantic keypoints $\boldsymbol{K}$ on the novel clothes. With the pixel locations of these keypoints, we further compute their 3D coordinates using the corresponding depth map and camera parameters.
Figure 5: Skill library. CLASP skill library consists of 5 basic skills: grasp, moveto, release, rotate and pull.
@@ 5.1 Skill Library
For general-purpose clothes manipulation, we build a skill library first. Instead of manually designing skills, we prompt an LLM to discover basic clothes manipulation skills. The LLM’s commonsense knowledge ensures these skills are general and sufficient for various clothes and manipulation tasks. Specifically, the LLM is prompted with a chain-of-thought strategy (Wei et al., 2022) to (i) generate examples of clothes manipulation tasks, (ii) break down these tasks into basic actions, and (iii) summarize the actions identified in step (iii) to derive a set of basic skills. In this way, we identify the basic skill set $\mathcal{A}$, including grasp: pick up a part of the clothes, moveto: transport the clothes to a target position, release: drop a part of the clothes, rotate: rotate the clothes, and pull: stretch the clothes for flattening. Figure 5 visualizes the basic skills discovered by the LLM. Since semantic keypoints serve as the effective affordances for clothes manipulation and define where clothes are more likely to be manipulated, we implement these basic skills as policies parameterized by contact point positions. Specifically, we define heuristic rules to create waypoints based on the contact point positions for each basic skill. Once the waypoints are determined, a motion planning algorithm is applied to generate the complete trajectory for the robot’s execution. Specifically, we implement each basic skill as follows:
@@
release()(): The skill release drops clothes parts by opening grippers. After the release action, the observation is updated, and CLASP extracts new semantic keypoints.
@@
pull(l)(l): The skill pull stretches the clothes to flatten them, where $l$ represents the stretch distance. In this paper, we set the stretch distance $l$ to 10% of the distance between the grippers of the two arms. A potential improvement could involve stretching the clothes as much as possible without tearing, monitored through a force sensor or a front-view camera.
@@ 5.2 Task Planning
Figure 6: Example of few-shot prompting for the VLM-powered task planner.
Algorithm 1 Task Planning Procedure
1:Instruction $\mathcal{L}$, image $I$, keypoints $\boldsymbol{K}$, skill set $\mathcal{A}$
2:Valid task plan or failure reason
3:if CheckCompletion($\mathcal{L},I,\boldsymbol{K}$) then
4:  return Task Completed
5:end if
@@
7:  $\boldsymbol{W}\leftarrow$ ProposeTaskPlan($\mathcal{L},I,\boldsymbol{K},\mathcal{A}$)
8:  $valid\leftarrow$ PlanVerification($\boldsymbol{W},\mathcal{A},\boldsymbol{K}$)
9:until validvalid or attempts are exhausted
10:return $\boldsymbol{W}$ or failure reason
Algorithm 2 Check validity of a proposed task plan $\boldsymbol{W}$
1:PlanVerification($\boldsymbol{W},\mathcal{A},\boldsymbol{K}$)
2:for each $\boldsymbol{w}_{j}=(a_{j},c_{j}^{1},c_{j}^{2})$ in $\boldsymbol{W}$ do
3:  if $a_{j}\notin\mathcal{A}$ or $c_{j}^{1,2}\notin\boldsymbol{K}$ then
4:   return False
5:  else if not MotionFeasibility($a_{j},c_{j}^{1},c_{j}^{2}$) then
6:   return False
7:  end if
8:end for
9:return True
After establishing the skill library for clothes manipulation, we leverage the VLM to perform task planning over semantic keypoints, utilizing its commonsense knowledge acquired from internet-scale data (Zhao et al., 2024). The task planning procedure (Algorithm 1) consists of three main modules: CheckCompletion, ProposeTaskPlan, and PlanVerification.
CheckCompletion: Given a language instruction $\mathcal{L}$, an observation image $I$, and extracted semantic keypoints $\boldsymbol{K}$, CLASP first determines whether the task specified by $\mathcal{L}$ has already been completed. If the completion condition is satisfied, the procedure terminates.
ProposeTaskPlan: If the task remains incomplete, the VLM generates a candidate task plan $\boldsymbol{W}=(\boldsymbol{w}_{1},\boldsymbol{w}_{2},\ldots,\boldsymbol{w}_{H})$. Each subtask $\boldsymbol{w}_{j}=(a_{j},c_{j}^{1},c_{j}^{2})$ consists of a basic skill $a_{j}$ and up to two contact points $(c_{j}^{1},c_{j}^{2})$ for the dual-arm manipulation system. The skill $a_{j}$ is selected from the predefined set $\mathcal{A}$ introduced in Sec. 5.1, while the contact points are chosen from the semantic keypoints $\boldsymbol{K}$. Semantic keypoints define where clothes can be manipulated and serve as contact point candidates in high-level task planning. To improve robustness, we employ in-context learning by providing the VLM with a few-shot prompt that includes example pairs of natural-language instructions, observation images annotated with semantic keypoints, and corresponding task plans (Figure 6). These examples define the expected input–output structure and demonstrate how to ground instructions into executable skills over semantic keypoints.
PlanVerification: The proposed candidate task plan $\boldsymbol{W}$ is then validated through the procedure described in Algorithm 2. For each subtask $\boldsymbol{w}_{j}$, CLASP first performs syntax and semantic checks—verifying that the skill $a_{j}$ exists in $\mathcal{A}$ and that contact points $c_{j}^{1,2}$ correspond to valid semantic keypoints. Next, it evaluates geometric feasibility via motion planning. A subtask passes verification if a feasible trajectory is found; otherwise, the failure reason is used to prompt VLM to generate a new task plan.
@@
Once a verified task plan is obtained, CLASP executes the subtasks sequentially. After each subtask, the RGB-D observation is updated. However, due to potential occlusion caused by the robot, semantic keypoints are re-extracted only after executing a release action. Following each keypoint update, the VLM is prompted to replan based on the updated observation and keypoints. This process repeats iteratively until the task is successfully completed.
@@ 6 Experiments
Figure 7: Qualitative results on semantic keypoint extraction. Our method can achieve open-category semantic keypoint extraction under irregular deformation and occlusion.
@@
We evaluate the performance of CLASP’s submodules: semantic keypoint extraction and task planning. Sec. 6.1 shows that the proposed method reliably extracts semantic keypoints across various clothes, even under occlusion and deformation. Sec. 6.2 shows that the task planner generates robust, executable plans, with closed-loop planning substantially improving performance on crumpled clothes.
@@ 6.1 Semantic Keypoint Extraction Experiments
Table 1: Semantic keypoint extraction performance.
@@
We first evaluated the performance of CLASP in extracting effective semantic keypoints. Since no ground-truth semantic keypoints are available, we provide annotated prototype images to three human experts and ask them to label semantic keypoints on images of novel clothes. In this way, we establish a testing set of 120 images from 15 different clothes. In these images, clothes have diverse configurations, poses, occlusions, and deformations. We then compare the semantic keypoints extracted by CLASP against those labeled by the experts.
The evaluation utilizes two standard metrics for keypoint extraction: Average Keypoint Distance (AKD) and Average Precision (AP). AKD calculates the average distance between the extracted keypoints and the ground-truth keypoints, while AP measures the proportion of correctly extracted keypoints under different thresholds. Since the resolution of the observed RGB-D images is 720×1280, we used thresholds of 15, 30, and 45 pixels for AP calculation.
@@
Considering the resolution and uncertainty in human expert labeling, such as identifying the center of clothes, the results can demonstrate that our proposed method performs well in extracting semantic keypoints. Figure 7 shows some qualitative results of semantic keypoint extraction. The experiment results demonstrate that our two-stage semantic keypoint extraction pipeline is effective in extracting semantic keypoints of clothes even under occlusion and deformation. The prior knowledge from the VLM and vision foundation models ensures the generalization to diverse clothes types. Besides, coarse-to-fine keypoint matching ensures spatial precision and makes semantic keypoint extraction robust. Our method enables effective semantic keypoint extraction on open-category clothes using a single unannotated prototype image per clothes category, without the need for large-scale labeled data.
Table 2: Task planning performance.